OpenStudy (anonymous):

please help me i do not get this at all.. statements and reasonng.. given:RS= 1/2 RT prove S is the midpoint of line RT

OpenStudy (anonymous):

so far i have STATEMENT REASONING 1) RS= 1/2RT given 2) 2RS=RT multiplication property 3) ? Segment Addition postulate 4) 2RS=RS+ST ? 5) RS+ST ? 6) ? Defintion of midpoint

OpenStudy (anonymous):

im thinking 6 is S is the midpoint of line RT

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

I also think that number 3 is RS+ST=RT

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

and i think number 5 is division property, so i just need someone to make sure all of it is correct and help me with number 4 please..